OpenStudy (anonymous):

Simplify. (4x^2-3xy-5y^2)-(2x^2+xy-3y^2)

OpenStudy (anonymous):

\[(4x ^{2}-3xy-5y ^{2})-(2x ^{2}+xy-3y ^{2)}\]

OpenStudy (anonymous):

.removing the parenthesis of th given eq. (4x^2-3xy-5y^2)-(2x^2+xy-3y^2) gives, = 4x^2-3xy-5y^2 - 2x^2 - xy + 3y^2 now try to simplify by adding and subtracting =?

OpenStudy (anonymous):

@Eric668668? can you do it ?

OpenStudy (anonymous):

It comes out to be \[2x ^{2}-4xy-8y ^{2}\]

OpenStudy (anonymous):

add or subtract liked terms on = 4x^2-3xy-5y^2 - 2x^2 - xy + 3y^2 or this is the same as = 4x^2 - 2x^2 - 3xy - xy - 5y^2 +3y^2 =?

OpenStudy (anonymous):

wolfram it!

OpenStudy (anonymous):

you almost got it except in this one - 5y^2 +3y^2

OpenStudy (anonymous):

(-5y^2)-(-3y^2)= - 5y^2 +3y^2

OpenStudy (anonymous):

Okay, this is what I got: \[2x ^{4}-2xy-8y ^{4}\]

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

on the first solution you got you almost got it except in this one - 5y^2 +3y^2=?

OpenStudy (anonymous):

it shouldnt be in the 4th power

OpenStudy (anonymous):

So the answer to - 5y^2 +3y^2=? is -16y^2

OpenStudy (anonymous):

hhmm its just -5+3 = +3-5 =__? its not 8 or 16

OpenStudy (anonymous):

So it's -2^2

OpenStudy (anonymous):

yeahhh you are correct its -2 y^2 .... :D

OpenStudy (anonymous):

So the full answer is 2x2−4xy−2y2

OpenStudy (anonymous):

yup you are correct ... :D

OpenStudy (anonymous):

Ohhh YAY!!! Thank You Soooooo Much!!!!!!!!!!!!!!!!!!!!!!!!!! :D

OpenStudy (anonymous):

ok good luck njow ....enjoy solving :D